2x+7y=5 -- [find two points on this line]
Answer by TimothyLamb(4379)

p1:
try x = -29
2x + 7y = 5
2(-29) + 7y = 5
-58 + 7y = 5
7y = 5 + 58
7y = 63
y = 9
p1 = (-29, 9)
---
p2:
try x = -15
2x + 7y = 5
2(-15) + 7y = 5
-30 + 7y = 5
7y = 5 + 30
7y = 35
y = 5
p2 = (-15, 5)
---
Answer:
Draw a line between the points p1 and p2 given above.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php